OpenStudy (anonymous):

How do i convert degree into radiant. -37° 30' ???? helP!

OpenStudy (anonymous):

360 degrees = 2*pi radians. 180 degrees = pi radians pi = 3.14

OpenStudy (anonymous):

i know that but the answer is coming -5pi/24 radiant how?

OpenStudy (anonymous):

Can you clarify your question: is it -30 degrees to radians? and 30 degrees to radians???

OpenStudy (anonymous):

it is -37° 30'

OpenStudy (anonymous):

-37° 30' + (1/2) as 1° = 60'................then we would take the LCM which is coming -74 +1/2 How??? it is then converted as -75/2 °

OpenStudy (anonymous):

Now convert it into radians..

OpenStudy (anonymous):

i am not getting how -74+1/2 can be -75/2 ??

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

How did you get -74 ??

OpenStudy (anonymous):

it is: -37 + 1/2 Now take the LCM you wiil get that..

OpenStudy (anonymous):

-74 +1/2

OpenStudy (anonymous):

30'/60 should = number of degrees in 30' add to -37. Then convert to radians. That's what makes sense to me. But I don't think we know enough about your question, and the problem.

OpenStudy (anonymous):

\[-37 + \frac{1}{2} \implies -\frac{37 \times 2 + 1}{2} \implies \frac{-75}{2}\]

OpenStudy (anonymous):

Now multiply it with pi/180..

OpenStudy (anonymous):

i thought that -74 + 1/2 = -73/2 ??

OpenStudy (anonymous):

\[\large \frac{-\cancel{75}^{5}}{2} \times \frac{\pi}{\cancel{180}^{12}} = \frac{-5 \pi}{24}\]

OpenStudy (anonymous):

No not is that..

OpenStudy (anonymous):

For just two minutes forget the negative sign.. Just consider it as +..

OpenStudy (anonymous):

It is basically : \(-(37.30')\)

OpenStudy (anonymous):

minus is just outside of everything..

OpenStudy (anonymous):

you mean i need to do this just in trignometry ??

OpenStudy (anonymous):

I show you: Remember - is not with 37... \[\large -(37.30') \implies -(37 + \frac{1}{2}) \implies -(\frac{75}{2})\]

OpenStudy (anonymous):

If you want to apply - sign then you will distribute it to 30' as well and not only to 37..